Saxbury Corporation's relevant range of activity is 3,000 units to 7,000 units. When it produces and sells 4,100 units, its average costs per unit are as follows: Average Cost per Unit Direct materials $ 6.70 Direct labor $ 3.40 Variable manufacturing overhead $ 1.50 Fixed manufacturing overhead $ 3.80 Fixed selling expense $ 0.70 Fixed administrative expense $ 0.55 Sales commissions $ 0.45 Variable administrative expense $ 0.45 Required:

a. For financial reporting purposes, what is the total amount of product costs incurred to make 4,100 units?

b. For financial reporting purposes, what is the total amount of period costs incurred to sell 4,100 units?

c. If 5,100 units are sold, what is the variable cost per unit sold? (Round "Per unit" answer to 2 decimal places.)

d. If 5,100 units are sold, what is the total amount of variable costs related to the units sold?

e. If 5,100 units are produced, what is the average fixed manufacturing cost per unit produced? (Round "Per unit" answer to 2 decimal places.)

f. If 5,100 units are produced, what is the total amount of fixed manufacturing cost incurred?

g. If 5,100 units are produced, what is the total amount of manufacturing overhead cost incurred? What is this total amount expressed on a per unit basis? (Round "Per unit" answer to 2 decimal places.)

h. If the selling price is $24.20 per unit, what is the contribution margin per unit sold? (Round "Per unit" answer to 2 decimal places.)

i. If 3,100 units are produced, what is the total amount of direct manufacturing cost incurred?

j. If 3,100 units are produced, what is the total amount of indirect manufacturing cost incurred?

k. What incremental manufacturing cost will the company incur if it increases production from 4,100 to 4,101 units? (Round "Per unit" answer to 2 decimal places.)
Answer #1

a. Total amount of product cost:

Direct material (4,100*$6.70) $27,470
Direct labor (4,100*$3.40) 13,940
Variable manufacturing overhead (4,100*$1.50) 6,150
Fixed manufacturing overhead (4,100*$3.80) 15,580
Total product cost $63,140

b. Total amount of period cost:

Fixed selling expense (4,100*$0.70) $2,870
Fixed administrative expense (4,100*$0.55) 2,255
Sales commission (4,100*$0.45) 1,845
Variable administrative expense (4,100*$0.45) 1,845
Total period costs $8,815

c. Variable cost per unit sold :

Direct material $6.70
Direct labor 3.40
Variable manufacturing overhead 1.50
Sales commission 0.45
Variable administrative expense 0.45
Variable cost of unit sold $12.5

d. Total amount of variable cost = 5,100*$12.5 = $63,750

e. Average fixed manufacturing cost per unit = 4,100*$3.80 / 5,100 = $3.05

f. Total amount of fixed manufacturing cost incurred = $15,580

g. Total amount of manufacturing overhead cost incurred = 5,100*$11.6 ($6.70+3.40+1.50) + 15,580

Total amount of manufacturing overhead cost incurred = $59,160+15,580 = $74,740

Per unit = $74,740 / 5,100 = $14.65 per unit

h.

Selling price per unit $24.20
Variable cost per unit -12.5
Contribution margin per unit sold $11.7

i. Total amount of direct manufacturing cost incurred :

Direct material (3,100*$6.70) $20,770
Direct labor (3,100*$3.40) 10,540
Direct manufacturing cost $31,310

j.   Total amount of indirect manufacturing cost incurred :

Variable manufacturing overhead (3,100*$1.50) $4,650
Fixed manufacturing overhead 15,580
Indirect manufacturing cost $20,230

k. Incremental manufacturing costs = $6.70+3.40+1.50 = $11.6